# ReCamMaster: Camera-Controlled Generative Rendering from A Single Video

### v1

@@ Abstract
Camera control has been actively studied in text or image conditioned video generation tasks. However, altering camera trajectories of a given video remains under-explored, despite its importance in the field of video creation. It is non-trivial due to the extra constraints of maintaining multiple-frame appearance and dynamic synchronization. To address this, we present ReCamMaster, a camera-controlled generative video re-rendering framework that reproduces the dynamic scene of an input video at novel camera trajectories. The core innovation lies in harnessing the generative capabilities of pre-trained text-to-video models through a simple yet powerful video conditioning mechanism—its capability often overlooked in current research. To overcome the scarcity of qualified training data, we construct a comprehensive multi-camera synchronized video dataset using Unreal Engine 5, which is carefully curated to follow real-world filming characteristics, covering diverse scenes and camera movements. It helps the model generalize to in-the-wild videos. Lastly, we further improve the robustness to diverse inputs through a meticulously designed training strategy. Extensive experiments tell that our method substantially outperforms existing state-of-the-art approaches and strong baselines. Our method also finds promising applications in video stabilization, super-resolution, and outpainting.
@@ 1 Introduction
To address these challenges, we present ReCamMaster, a camera-controlled generative video re-rendering framework that can regenerate in-the-wild videos at novel camera trajectories. The core innovation is to utilize the generative capabilities of pre-trained text-to-video models through an elegant yet powerful video conditioning mechanism. Compared to alternatives, our method shows superior performance in preserving both the visual and dynamic characteristics of the input video. Notably, this conditioning mechanism indeed exhibits remarkable potential as a versatile solution for conditional generation tasks, which is overlooked in current research. Since no qualified training data are publicly available, we develop a large-scale multi-camera synchronized video dataset using Unreal Engine 5, which contains 136K realistic videos shooting from 13.6K different dynamic scenes in 40 high-quality 3D environments with 122K different camera trajectories. Particularly, it is carefully curated to simulate real-world filming characteristics, which is proved to bring advantage to striking domain alignment to in-the-wild videos. Lastly, we further improve the robustness of our model to diverse inputs through a meticulously designed training strategy.
@@
  • We introduce a high-quality multi-camera synchronized video dataset featuring diverse camera trajectories. This dataset will be publicly released to advance research in camera-controlled video generation, 4D reconstruction, and related fields.
@@ 2 Related Works
With the success of text-to-video generation models [6, 7, 60, 34, 10], the introduction of other conditional signals for controllable video generation has been widely studied [61, 16, 54, 11]. In camera-controlled video generation [59, 2, 64, 65], researchers aim to incorporate camera parameters into video generation models to control the viewpoint of the output video. AnimateDiff [17] introduces various motion LoRAs [21] to learn specific patterns of camera movements. MotionCtrl [49] encodes 6DoF camera extrinsics and injects them into the diffusion model, fine-tuning on video-camera pair data to achieve video generation with arbitrary trajectory control. CameraCtrl [18] further improves the accuracy and generalizability of single-sequence camera control with a dedicatedly designed camera encoder. CVD [28] achieve multi-sequence camera control with the proposed cross-video synchronization module. AC3D [1] conducts an in-depth investigation into camera motion knowledge within diffusion transformers, achieving camera-controlled generation with enhanced visual quality. Meanwhile, training-free methods have also been explored [20, 22, 30].
@@
Video-to-video generation [46, 47, 33] is explored in various tasks, including video editing [58, 32], video outpainting [44, 8], video super-resolution [66, 56], etc. Most relevant to our work is synthesizing novel viewpoint videos based on a given video and camera parameters [43, 62, 5, 15, 38]. GCD [43] pioneered camera-controlled video-to-video generation by training a video generation model with paired video data created by the Kubric simulator [14]. Although it performs well on in-domain data, GCD’s generalization capability is limited by the significant domain gap between the training data and real-world videos. Recapture [62] propose a practical solution with per-video optimization. Several concurrent works, including DaS [15], Trajectory-Attention [52], and GS-DiT [5], achieve 4D-consistent video-to-video generation by extracting dynamic information from the original video using 3D point tracking [26, 51] and incorporate it as a condition into the video generator. This paradigm is promising when synchronized multi-view videos are unavailable. Nevertheless, the generation quality is limited by the accuracy of the point tracking methods. In the field of 4D object generation, video-to-video generation can be achieved by 1) training a multi-view video generator [53, 50], or 2) following a reconstruct-and-render pipeline via 4D reconstruction methods [63, 29]. In this paper, we focus on open-domain camera-controlled video generation. To achieve this, we first create a large multi-camera synchronized dataset with diverse camera trajectories. We also propose a novel video conditioning method equipped with improved training strategy to achieve better 4D consistency and generation ability.
@@ 3 Multi-Cam Video: A High-Quality Multi-Camera Synchronized Video Dataset
We build the entire data rendering pipeline in Unreal Engine 5 [12]. Specifically, we first collect multiple 3D environments as “backgrounds”. Then, we place animated characters within these environments as the “main subjects” of the videos. We then position multiple cameras facing the subjects and moving along predefined trajectories to simulate the process of simultaneous shooting. This allows us to render datasets with synchronized cameras that include dynamic objects. To scale up the data volume, we construct a set of camera movement rules for automatically batch generation of natural and diverse camera trajectories. Additionally, we randomly combined different characters and actions across different video sets. In total, we obtain 136K visually-realistic videos shooting from 13.6K different dynamic scenes in 40 high-quality 3D environments with 122K different camera trajectories. Fig. 8 in Appendix shows some of the rendered video frames. In Appendix C.1, we conduct an ablation study comparing model performance when trained on a dataset with limited scenes and camera trajectories versus our comprehensive dataset. This highlights the significance of our high-quality dataset. For more details about the dataset, please refer to Appendix B.
@@ 4 Camera-Controlled Video Re-Generation
Given a source video $V_{s}\in\mathbb{R}^{f\times c\times h\times w}$, we aim to synthesize a target video $V_{t}\in\mathbb{R}^{f\times c\times h\times w}$ sharing the same dynamic scene but presenting at specified camera trajectories denoted by $\texttt{cam}_{t}\coloneqq[{R},{t}]\in\mathbb{R}^{f\times 3\times 4}$. Particularly, $V_{t}$ should conform to the multiple-frame appearance and synchronized dynamics of $V_{s}$. To achieve this, we propose to harness the generative capability of pre-trained text-to-video diffusion models [60, 7] by imposing dual conditions, i.e. the source video and target camera trajectories through a meticulously designed framework. The overview of the model is depicted in Fig. 3.
@@ 4.2 Conditional Video Injection Mechanism
Recent attempts [43, 5] incorporate the source video by concentrating the latent of the source video $x_{s}$ with the noised latent of the target video $x_{t}$ along the channel dimension, and add additional input channels to the input convolutional layer. To be specific, a variational autoencoder with encoder $\mathcal{E}$ is utilized to project the source video $V_{s}$ and target video $V_{t}$ to the latent space, $z_{s}=\mathcal{E}(V_{s}),z_{t}=\mathcal{E}(V_{t})$, where $z_{s},z_{t}\in\mathbb{R}^{b\times f\times c\times h\times w}$ are the latent of source video and target video with $f$ frames, $c$ channels, and spatial size of $h\times w$. Then, the source latent is concentrated with the noised target latent along the channel dimension, and pathify into $f\times h\times w$ tokens with several convolutional layers:
@@
where $x_{t}\in\mathbb{R}^{b\times f\times s\times d}$, $s=h\times w$, $d$ is the channel dimension for the latent diffusion model. We denote this conditioning technique as the “channel-dimension conditioning” and illustrated in Fig. 3(b).
@@
To achieve multi-view video generation, [3] propose to introduce a plug-and-play module for feature aggregation across views. Given a source video latent $z_{s}=\mathcal{E}(V_{s})$ and a target video latent $z_{t}=\mathcal{E}(V_{t})$, and patchify and forward to the diffusion transformer respectively:
@@
where $x_{i}\in\mathbb{R}^{b\times 2f\times s\times d}$ is the input of diffusion transformer. In other words, the input token number is doubled compared to the vanilla text-to-video generation process. Moreover, we do not introduce additional attention layers for feature aggregation across the source video and the target video, since self-attention is performed between all tokens in the 3D (spatial-temporal) attention layers. We denote this conditioning technique as the “frame-dimension conditioning” and illustrated in Fig. 3(a).
@@ 5.1 Experiment Settings
Evaluation Metrics. We mainly evaluate the proposed method in terms of camera accuracy, source-target synchronization, and visual quality. For camera accuracy, we use GLOMAP [35] to extract the camera pose sequence of the generated videos, and calculate the rotation error and translation error, denoted as RotErr and TransErr respectively [18]. In terms of synchronization, we utilize the state-of-the-art image matching method GIM [40] to calculate the number of matching pixels with confidence greater than the threshold, denoted as Mat. Pix.. Furthermore, we calculate the FVD-V score in SV4D [53], and the average CLIP similarity between source and target frames at the same timestamp, denoted as CLIP-V [28]. For visual quality, we divide it into fidelity, coherence with text, and temporal consistency, and quantified them with Fréchet Image Distance [19] (FID) and Fréchet Video Distance [42] (FVD), CLIP-T, and CLIP-F respectively. CLIP-T refers to the average CLIP [37] similarity of each frame and its corresponding text prompt, and CLIP-F is the average CLIP similarity of adjacent frames. We also evaluate our method on the widely-used VBench [23] metrics.
@@ 5.3 More Analysis and Ablation Studies
Figure 5: Ablation on video conditioning techniques. We compared the channel-/view- concatenation schemes proposed by previous methods and frame-concatenation in ReCamMaster. We observed that both channel-conditioning and view-conditioning schemes suffer from significant artifacts, content inconsistency and asynchronous dynamics with respect to the original video.
In Sec. 4.2, we introduce a novel video conditioning scheme that concatenates the tokens of a source video with the target video tokens along the frame dimension. To verify its effectiveness, we compare our “frame concatenation” technique with the “channel concatenation” used in baseline methods [43, 5] and the “view concatenation” in [3]. We performed both qualitative and quantitative comparisons, as shown in Fig. 5 and Tab. 3. We only altered the video conditioning method while keeping the camera injection method, training data, and training steps unchanged. The results clearly show that our conditioning technique significantly enhances the model’s performance. In the second and third rows of Fig. 5, the generated videos exhibit the rendering style of the training dataset, and the hand movements become unsynchronized with the original video. In contrast, our method preserves the person’s identity and maintains synchronization even during rapid and complex movements.
@@
In Section 4.4, we enhance ReCamMaster’s robustness with improved training strategies. We fine-tune only the spatial-temporal (3D) attention layers and freeze other parameters. Additionally, we drop $f$ frames and $f-1$ frames latent of the source video with a 20% probability to improve generation capability. Table 4 shows the improvement using visual quality metrics. The ‘baseline’ is vanilla training with reconstruction loss. Results indicate that both techniques improve visual quality, with their combination yielding the best performance. Dropping the source latent during training also allows our method to support camera-controlled T2V, I2V, and V2V generation simultaneously, as shown in Fig. 9.
@@ 6 Applications of ReCamMaster
Video Stabilization. It is usually not easy to obtain stable videos when recording video while moving. Video stabilization techniques [48] aim to smooth out camera movements to produce easy-to-watch videos, which can be achieved by adjusting the camera trajectories via ReCamMaster. We evaluate our methodd using unsteady videos from the DeepStab [45] dataset. It can be observed that the model stabilizes the video while preserving the scenes and actions from the original video. Best view the video results on the project page.
Video Super-Resolution. Thanks to the generation capability of the diffusion model, we can input zoom-in camera trajectories into ReCamMaster to achieve video local super-resolution, as more details are observed in Fig. 6.
@@ Appendix B Details of Data Construction
In this section, we provide the detailed description of the rendered dataset used to train ReCamMaster.
@@
We collect 40 different 3D environments assets from https://www.fab.com/. To minimize the domain gap between rendered data and real-world videos, we primarily select visually realistic 3D scenes, while choosing a few stylized or surreal 3D scenes as a supplement. To ensure data diversity, the selected scenes cover a variety of indoor and outdoor settings, such as city streets, shopping malls, cafes, office rooms, and countryside.
@@
We chose two set of commonly used camera parameters: focal=35mm, aperture=2.8 and focal=24mm, aperture=10.
@@ C.1 Ablation on Dataset Construction
In our experiments, we find that constructing a diverse training dataset that closely resembles the distribution of real-world videos significantly enhances the model’s generalization ability. To demonstrate this, we quantitatively compared the model performance trained on the “toy data” constructed in the early stages of the experiment and the “high-quality data” used in this paper. Specifically, the “toy data” was constructed with 500 scenes in a single 3D environment, and we manually created 20 camera trajectories assigned to 5000 cameras. As a result, this dataset lacked diversity in terms of scenes and camera trajectories, which limited the model’s generalization ability on in-the-wild videos and novel camera trajectories. In contrast, the “high-quality data” used in the paper contains 136K videos shot from 13.6K different dynamic scenes in 40 3D environments with 122K different camera trajectories. We present the results in Tab. 5. It is observed that the model shows significant improvements in visual quality, camera accuracy, and synchronization metrics.

### v2

@@ Abstract
Camera control has been actively studied in text or image conditioned video generation tasks. However, altering camera trajectories of a given video remains under-explored, despite its importance in the field of video creation. It is non-trivial due to the extra constraints of maintaining multiple-frame appearance and dynamic synchronization. To address this, we present ReCamMaster, a camera-controlled generative video re-rendering framework that reproduces the dynamic scene of an input video at novel camera trajectories. The core innovation lies in harnessing the generative capabilities of pre-trained text-to-video models through a simple yet powerful video conditioning mechanism—its capability is often overlooked in current research. To overcome the scarcity of qualified training data, we construct a comprehensive multi-camera synchronized video dataset using Unreal Engine 5, which is carefully curated to follow real-world filming characteristics, covering diverse scenes and camera movements. It helps the model generalize to in-the-wild videos. Lastly, we further improve the robustness to diverse inputs through a meticulously designed training strategy. Extensive experiments show that our method substantially outperforms existing state-of-the-art approaches. Our method also finds promising applications in video stabilization, super-resolution, and outpainting. Our code and dataset are publicly available at: https://github.com/KwaiVGI/ReCamMaster.
@@ 1 Introduction
To address these challenges, we present ReCamMaster, a camera-controlled generative video re-rendering framework that can regenerate in-the-wild videos at novel camera trajectories. The core innovation is to utilize the generative capabilities of pre-trained text-to-video models through an elegant yet powerful video conditioning mechanism. Compared to alternatives, our method shows superior performance in preserving both the visual and dynamic characteristics of the input video. Notably, this conditioning mechanism indeed exhibits remarkable potential as a versatile solution for conditional generation tasks, which is overlooked in current research. Since no qualified training data are publicly available, we develop a large-scale multi-camera synchronized video dataset using Unreal Engine 5, which contains 136K realistic videos shot from 13.6K different dynamic scenes in 40 high-quality 3D environments with 122K different camera trajectories. Particularly, it is carefully curated to simulate real-world filming characteristics, which is proven to bring an advantage to striking domain alignment to in-the-wild videos. Lastly, we further improve the robustness of our model to diverse inputs through a meticulously designed training strategy.
@@
  • We introduce a high-quality multi-camera synchronized video dataset featuring diverse camera trajectories. This dataset has been publicly released to advance research in camera-controlled video generation, 4D reconstruction, and related fields.
@@ 2 Related Works
With the success of text-to-video generation models [6, 7, 62, 35, 10], the introduction of other conditional signals for controllable video generation has been widely studied [63, 16, 56, 11]. In camera-controlled video generation [61, 2, 67, 68], researchers aim to incorporate camera parameters into video generation models to control the viewpoint of the output video. AnimateDiff [17] introduces various motion LoRAs [21] to learn specific patterns of camera movements. MotionCtrl [51] encodes 6DoF camera extrinsics and injects them into the diffusion model, fine-tuning on video-camera pair data to achieve video generation with arbitrary trajectory control. CameraCtrl [18] further improves the accuracy and generalizability of single-sequence camera control with a dedicatedly designed camera encoder. CVD [28] achieves multi-sequence camera control with the proposed cross-video synchronization module. AC3D [1] conducts an in-depth investigation into camera motion knowledge within diffusion transformers, achieving camera-controlled generation with enhanced visual quality. Meanwhile, training-free methods have also been explored [20, 22, 30].
@@
Video-to-video generation [48, 49, 34] is explored in various tasks, including video editing [60, 33], video outpainting [46, 8], video super-resolution [69, 58], etc. Most relevant to our work is synthesizing novel viewpoint videos based on a given video and camera parameters [45, 65, 5, 15, 39]. GCD [45] pioneered camera-controlled video-to-video generation by training a video generation model with paired video data created by the Kubric simulator [14]. Although it performs well on in-domain data, GCD’s generalization capability is limited by the significant domain gap between the training data and real-world videos. Recapture [65] proposes a practical solution with per-video optimization. Several concurrent works [64, 39, 15, 54, 5], achieve 4D-consistent video-to-video generation by extracting dynamic information from the original video using 3D point tracking [26, 53] and incorporating it as a condition into the video generator. This paradigm is promising when synchronized multi-view videos are unavailable. Nevertheless, the generation quality is limited by the accuracy of the point-tracking methods. In the field of 4D object generation, video-to-video generation can be achieved by 1) training a multi-view video generator [55, 52], or 2) following a reconstruct-and-render pipeline via 4D reconstruction methods [66, 29]. [32, 42] achieve 3D/4D scene reconstruction with videos generated with diffusion models. In this paper, we focus on open-domain camera-controlled video generation. To achieve this, we first create a large multi-camera synchronized dataset with diverse camera trajectories. We also propose a novel video conditioning method to achieve better generation ability.
@@ 3 Multi-Cam Video: A High-Quality Multi-Camera Synchronized Video Dataset
We build the entire data rendering pipeline in Unreal Engine 5 [12]. Specifically, we first collect multiple 3D environments as “backgrounds". Then, we place animated characters within these environments as the “main subjects" of the videos. We then position multiple cameras facing the subjects and moving along predefined trajectories to simulate the process of simultaneous shooting. This allows us to render datasets with synchronized cameras that include dynamic objects. To scale up the data volume, we construct a set of camera movement rules for automatically batch generation of natural and diverse camera trajectories. Additionally, we randomly combined different characters and actions across different video sets. In total, we obtain 136K visually-realistic videos shot from 13.6K different dynamic scenes in 40 high-quality 3D environments with 122K different camera trajectories. Fig. 8 in the Appendix shows some of the rendered video frames. In Appendix C.1, we conduct an ablation study comparing model performance when trained on a dataset with limited scenes and camera trajectories versus our comprehensive dataset. This highlights the significance of our high-quality dataset. For more details about the dataset, please refer to Appendix B.
@@ 4 Camera-Controlled Video Re-Generation
Given a source video $V_{s}\in\mathbb{R}^{f\times c\times h\times w}$, we aim to synthesize a target video $V_{t}\in\mathbb{R}^{f\times c\times h\times w}$ sharing the same dynamic scene but presenting at specified camera trajectories denoted by $\texttt{cam}_{t}\coloneqq[{R},{t}]\in\mathbb{R}^{f\times 3\times 4}$. Particularly, $V_{t}$ should conform to the multiple-frame appearance and synchronized dynamics of $V_{s}$. To achieve this, we propose to harness the generative capability of pre-trained text-to-video diffusion models [62, 7] by imposing dual conditions, i.e., the source video and target camera trajectories through a meticulously designed framework. The overview of the model is depicted in Fig. 3.
@@ 4.2 Conditional Video Injection Mechanism
Recent attempts [45, 5] incorporate the source video by concatenating the latent of the source video $x_{s}$ with the noised latent of the target video $x_{t}$ along the channel dimension, and add additional input channels to the input convolutional layer. To be specific, a variational autoencoder with encoder $\mathcal{E}$ is utilized to project the source video $V_{s}$ and target video $V_{t}$ to the latent space, $z_{s}=\mathcal{E}(V_{s}),z_{t}=\mathcal{E}(V_{t})$, where $z_{s},z_{t}\in\mathbb{R}^{b\times f\times c\times h\times w}$ are the latent of source video and target video with $f$ frames, $c$ channels, and spatial size of $h\times w$. Then, the source latent is concentrated with the noised target latent along the channel dimension, and patchified into $f\times h\times w$ tokens with several convolutional layers:
@@
where $x_{t}\in\mathbb{R}^{b\times f\times s\times d}$, $s=h\times w$, $d$ is the channel dimension for the latent diffusion model. We denote this conditioning technique as the “channel-dimension conditioning” and illustrate in Fig. 3(b).
@@
To achieve multi-view video generation, [3] proposes to introduce a plug-and-play module for feature aggregation across views. Given a source video latent $z_{s}=\mathcal{E}(V_{s})$ and a target video latent $z_{t}=\mathcal{E}(V_{t})$, and patchify and forward to the diffusion transformer respectively:
@@
where $x_{i}\in\mathbb{R}^{b\times 2f\times s\times d}$ is the input of diffusion transformer. In other words, the input token number is doubled compared to the vanilla text-to-video generation process. Moreover, we do not introduce additional attention layers for feature aggregation across the source video and the target video, since self-attention is performed between all tokens in the 3D (spatial-temporal) attention layers. We denote this conditioning technique as the “frame-dimension conditioning” and illustrate in Fig. 3(a).
@@ 5.1 Experiment Settings
Evaluation Metrics. We mainly evaluate the proposed method in terms of camera accuracy, source-target synchronization, and visual quality. For camera accuracy, we use GLOMAP [36] to extract the camera pose sequence of the generated videos, and calculate the rotation error and translation error, denoted as RotErr and TransErr respectively [18]. In terms of synchronization, we utilize the state-of-the-art image matching method GIM [41] to calculate the number of matching pixels with confidence greater than the threshold, denoted as Mat. Pix.. Furthermore, we calculate the FVD-V score in SV4D [55], and the average CLIP similarity between source and target frames at the same timestamp, denoted as CLIP-V [28]. For visual quality, we divide it into fidelity, coherence with text, and temporal consistency, and quantify them with Fréchet Image Distance [19] (FID) and Fréchet Video Distance [44] (FVD), CLIP-T, and CLIP-F, respectively. CLIP-T refers to the average CLIP [38] similarity of each frame and its corresponding text prompt, and CLIP-F is the average CLIP similarity of adjacent frames. We also evaluate our method on the widely used VBench [23] metrics.
@@ 5.3 More Analysis and Ablation Studies
Figure 5: Ablation on video conditioning techniques. We compared the channel-/view- concatenation schemes proposed by previous methods and frame-concatenation in ReCamMaster. We observed that both channel-conditioning and view-conditioning schemes suffer from significant artifacts, content inconsistency, and asynchronous dynamics with respect to the original video.
In Sec. 4.2, we introduce a novel video conditioning scheme that concatenates the tokens of a source video with the target video tokens along the frame dimension. To verify its effectiveness, we compare our “frame concatenation" technique with the “channel concatenation" used in baseline methods [45, 5] and the “view concatenation" in [3]. We performed both qualitative and quantitative comparisons, as shown in Fig. 5 and Tab. 3. We only altered the video conditioning method while keeping the camera injection method, training data, and training steps unchanged. The results clearly show that our conditioning technique significantly enhances the model’s performance. In the second and third rows of Fig. 5, the generated videos exhibit the rendering style of the training dataset, and the hand movements become unsynchronized with the original video. In contrast, our method preserves the person’s identity and maintains synchronization even during rapid and complex movements.
@@
In Section 4.4, we enhance ReCamMaster’s robustness with improved training strategies. We fine-tune only the spatial-temporal (3D) attention layers and freeze other parameters. Additionally, we drop $f$ frames and $f-1$ frames of the source video with a 20% probability to improve generation capability. Table 4 shows the improvement using visual quality metrics. The ‘baseline’ is vanilla training with reconstruction loss. Results indicate that all techniques improve visual quality, with their combination yielding the best performance. Dropping the source latent during training also allows our method to support camera-controlled T2V, I2V, and V2V generation simultaneously, as shown in Fig. 9.
@@ 6 Applications of ReCamMaster
Video Stabilization. It is usually not easy to obtain stable videos when recording video while moving. Video stabilization techniques [50] aim to smooth out camera movements to produce easy-to-watch videos, which can be achieved by adjusting the camera trajectories via ReCamMaster. We evaluate our method using unsteady videos from the DeepStab [47] dataset. It can be observed that the model stabilizes the video while preserving the content from the original video. Best viewed on the project page.
Video Super-Resolution. [^1] It’s not entirely equivalent to the super-resolution task, as ReCamMaster only enhances the resolution of patches in the central region. Thanks to the generation capability of the diffusion model, we can input ‘zoom-in’ camera trajectories into ReCamMaster to achieve video local super-resolution, as more details are observed in Fig. 6.
@@ Appendix B Details of Data Construction
In this section, we provide a detailed description of the rendered dataset used to train ReCamMaster.
@@
We collect 40 different 3D environments assets from https://www.fab.com/. To minimize the domain gap between rendered data and real-world videos, we primarily select visually realistic 3D scenes, while choosing a few stylized or surreal 3D scenes as a supplement. To ensure data diversity, the selected scenes cover a variety of indoor and outdoor settings, such as city streets, shopping malls, cafes, office rooms, and the countryside.
@@
We chose two sets of commonly used camera parameters: focal=35mm, aperture=2.8, and focal=24mm, aperture=10.
@@ C.1 Ablation on Dataset Construction
In our experiments, we find that constructing a diverse training dataset that closely resembles the distribution of real-world videos significantly enhances the model’s generalization ability. To demonstrate this, we quantitatively compared the model performance trained on the “toy data" constructed in the early stages of the experiment and the “high-quality data" used in this paper. Specifically, the “toy data" was constructed with 500 scenes in a single 3D environment, and we manually created 20 camera trajectories assigned to 5000 cameras. As a result, this dataset lacked diversity in terms of scenes and camera trajectories, which limited the model’s generalization ability on in-the-wild videos and novel camera trajectories. In contrast, the “high-quality data" used in the paper contains 136K videos shot from 13.6K different dynamic scenes in 40 3D environments with 122K different camera trajectories. We present the results in Tab. 5. It is observed that the model shows significant improvements in visual quality, camera accuracy, and synchronization metrics.